### v2

# MyShadeQuoteStyleleftmargin=0pt, rightmargin=0pt, backgroundcolor=gray!25, linewidth=0pt, skipbelow=skipabove=

Abstract.
Detecting vulnerabilities is vital for software security, yet deep learning-based vulnerability detectors (DLVD) face a data shortage, which limits their effectiveness. Data augmentation can potentially alleviate the data shortage, but augmenting vulnerable code is challenging and requires a generative solution that maintains vulnerability. Previous works have only focused on generating samples that contain single statements or specific types of vulnerabilities. Recently, large language models (LLMs) have been used to solve various code generation and comprehension tasks with inspiring results, especially when fused with retrieval augmented generation (RAG). Therefore, we propose VulScribeR, a novel LLM-based solution that leverages carefully curated prompt templates to augment vulnerable datasets. More specifically, we explore three strategies to augment both single and multi-statement vulnerabilities, with LLMs, namely Mutation, Injection, and Extension. Our extensive evaluation across three vulnerability datasets and DLVD models, using two LLMs, show that our approach beats two SOTA methods Vulgen and VGX, and Random Oversampling (ROS) by 27.48%, 27.93%, and 15.41% in f1-score with 5K generated vulnerable samples on average, and 53.84%, 54.10%, 69.90%, and 40.93% with 15K generated vulnerable samples. Our approach demonstrates its feasibility for large-scale data augmentation by generating 1K samples at as cheap as US\$ 1.88.
Vulnerability Augmentation, Deep Learning, Vulnerability Generation, Program Generation, Vulnerability Injection
ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software reliabilityccs: Software and its engineering Software safetyccs: Software and its engineering Parsersccs: Software and its engineering Software maintenance toolsccs: Security and privacy Vulnerability scannersccs: Security and privacy Software security engineeringccs: Software and its engineering Source code generation
1. Introduction
Vulnerability detection is a crucial task in software engineering. Recent Deep Learning-based Vulnerability Detection (DLVD) models (Li et al., 2018; Fu and Tantithamthavorn, 2022; Zhou et al., 2019; Chakraborty et al., 2020; Li et al., 2021; Wang et al., 2023b; Lu et al., 2024) have drawn more attention from research and industry communities due to their promising performance. However, these models suffer greatly from 1) a lack of sizable datasets and 2) discrepancies between the distribution of training and testing datasets. Data Sampling techniques (Yang et al., 2023; Ganz et al., 2023) can alleviate the former problem to an extent by upsampling vulnerable samples to balance the dataset and optionally adding more clean samples, which are abundant. However, Data Augmentation presents a more promising solution as it can help with both problems by generating large diverse datasets. For instance, Zhang (Zhang, 2021) and Liu et al. (Liu et al., 2024) focus on generating specific type(s) of bug/vulnerability. Similarly, Nong et al. proposed VulGen (Nong et al., 2023) and VGX (Nong et al., 2024) to mine single-statement vulnerability patterns and inject them into clean samples to generate vulnerable samples. Despite their contributions, these methods have limitations. They either generate only specific types of vulnerabilities (Zhang, 2021; Nong et al., 2024, 2023) or require sizable datasets to learn models for generating vulnerable code or pinpoint the location to inject the vulnerability (Nong et al., 2022, 2024, 2023), making them impractical for adoption in a real-world scenario.
Recently, Large Language Models (LLMs) have demonstrated promising results in code-related tasks, such as code understanding (White et al., 2024; Leinonen et al., 2023), generation (Gu, 2023), and vulnerability understanding (Fang et al., 2024; Islam et al., 2024). Intuitively, vulnerable data augmentation is a task that requires the ability of code comprehension, typically in vulnerability understanding, and code generation. LLMs’ strong ability for code comprehension and code generation fits this task well.
Therefore, to tackle both of those problems and overcome the limitations of previous works, we propose VulScribeR, a novel LLM-based solution that leverages carefully curated prompt templates to augment vulnerable datasets. More specifically, we design three strategies to generate vulnerable new code samples, namely Mutation, Injection, and Extension. Mutation refers to changing vulnerable code using code transformation that preserves the code’s semantics and syntactical correctness by prompting LLMs. Injection prompts LLMs to inject vulnerable segments of the vulnerable samples into a clean sample to create a new vulnerable sample. Extention aims to extend the vulnerable sample by adding parts of the logic of the clean sample to the vulnerable sample to improve the diversity of the context where the vulnerability could occur.
To generate realistic code, we employ Retrieval-Augmented Generation (RAG) in Injection and Extension. More specifically, in Injection, we retrieve vulnerable samples that are similar to the clean sample into which the vulnerable segment will be injected. Additionally, to enhance the diversity of the generated vulnerable samples, we employ a clustering process to ensure the data in all clusters have the choice to be retrieved during RAG. We employ RAG and clustering in Extension similarly.
Eventually, we employ a fuzzy parser as a filtering mechanism to discard invalid responses such as empty code samples and those with severe syntactical errors, to reduce noise in the generated samples.
To assess VulScribeR, we generated vulnerable samples for each strategy using two different LLMs, and evaluated the effectiveness of each data augmentation approach on three SOTA DLVD models (including both token-based and graph-based models) and three commonly used vulnerable datasets. We conducted various experiments in the form of the following Research Questions (RQs):
  • RQ1: How effective is VulScribeR compared to SOTA approaches?
    Results: Both Injection and Extension outperform the baselines and the Mutation by a large margin, while Injection provides a slightly higher performance compared to the Extension. For instance, Injection outperforms NoAug, Vulgen, VGX, and ROS by 30.80%, 27.48%, 27.93%, and 15.41% on average F1-score.
  • RQ2: How does RAG contribute to VulScribeR?
    Results: RAG component makes a significant contribution to Injection and Extension strategies.
  • RQ3: How does the quantity of the generated samples impact the effectiveness of vulnerability detection models?
    Results: Augmenting more vulnerable data by using Injection helps improve the effectiveness of DLVD models, while VulGen, VGX, and random over-sampling fail to improve the performance of DLVD models by augmenting more than 5K vulnerable samples. Our LLM-based approach is more feasible for large-scale vulnerable data augmentation.
In summary, our contributions are as follows:
  • To the best of our knowledge, we are the first to explore vulnerability augmentation using LLMs. We carefully designed three novel prompt templates with different strategies and proposed a comprehensive pipeline for vulnerability augmentation that can be used for large-scale vulnerability augmentation (that is as cheap as US\$1.88 per 1K samples).
  • We performed an extensive evaluation using two different LLMs, three DLVD models, and three datasets, demonstrating the superiority of VulScribeR compared to SOTA baselines, including the best latest techniques.
  • We made our source code, experimental results, and the augmented datasets publicly available for future research (dat, 2024).
2. Background & Related work
2.1. Vulnerability Augmentation
Vulnerable code samples are scarce in practice, and various approaches have been developed  (Nong et al., 2022, 2023, 2024; Liu et al., 2024; Ganz et al., 2023) to enhance the capability of vulnerability detection models by generating such data and expanding the datasets. For instance, Nong et al. (Nong et al., 2022) explored the feasibility of vulnerability injection through a neural code editing model (Yao et al., 2021), which is a DL-based model trained to transform clean code into vulnerable code via introducing a set of predicted changes to the AST of the program. This approach requires high-quality datasets and hence suffers from a chicken-egg dilemma and has limited use. Ganz et al.  (Ganz et al., 2023) present CodeGraphSMOTE, a method to generate new vulnerable samples by porting SMOTE (Chawla et al., 2002) to the graph domain. They convert code to graphs and use graph autoencoders to encode the graphs into their latent space. Then, it applies SMOTE (i.e. an interpolation-based sampling method) on the latent space to generate new vulnerable items.
Nong et al. later proposed VulGen (Nong et al., 2023), which mines vulnerabilities to collect single-statement vulnerable patterns. Then uses a modified transformer model to locate where injection should take place. VGX (Nong et al., 2024) is an improved version of VulGen where a significantly bigger vulnerability dataset was used to cover a wider range of single-statement vulnerabilities in the mining phase and the localization model was replaced with a semantics-aware contextualization transformer to predict the injection contexts better. However, both VulGen and VGX only target single statement vulnerabilities and require a comprehensive pattern mining process. Single statement vulnerabilities only account for less than 40% of all vulnerabilities in the Bigvul (Fan et al., 2020) dataset.
2.2. Source Code Augmentation with RAG and LLMs
Leveraging LLMs for source code augmentation is getting popular and it has been used for data augmentation for Semantic Code Search (Wang et al., 2023a) and Code Generation (Chen et al., 2024), both of which heavily rely on RAG. Wang et al. (Wang et al., 2023a) leverages RAG to retrieve similar $query-code$ pairs, then for each pair they instruct ChatGPT (OpenAI, 2022) to change the query and code in separate prompts using predefined rules included in the prompts. Then, they filter the augmented pairs using an encoder model that calculates their similarity score, removing the low-scoring pairs. Chen et al. (Chen et al., 2024) utilized Code Search to augment data for Code Generation. Specifically, they retrieve $context-function$ pairs, where the context can be either a function header or a comment, to populate prompt templates that instruct the LLM to generate code based on the given context and retrieved function. Their architecture uses three components, namely Retriever, Formulator, and Generator. Similar to previous studies, we also employ RAG to retrieve relevant context to enhance the code generation, while different from them, we focus on augmenting vulnerabilities and we employ a clustering process to ensure the diversity of generated data.
3. Methodology
Figure 1. The components of VulScribeR. All of the proposed strategies share the same generation and verifier components. The Formulator component of the mutation strategy only requires the vulnerable samples and receives input directly. (M, I, E) tags distinguish the difference between the input data flows for Mutation, Injection, and Extension strategies, respectively.
In this section, we elaborate on the details of our methodology, VulScribeR. We propose a RAG-based solution that leverages carefully curated prompt templates with a lenient filtering mechanism to generate vulnerable code snippets by utilizing LLMs, which allows us to generate diverse and realistic vulnerable code snippets.
We specifically design three prompting strategies to generate vulnerable code samples for data augmentation.
Consequently, We carefully designed 3 prompt templates for each augmentation strategy, namely Mutation, Injection, and Extension templates, on which we will elaborate in Section 3.1.
Figure 1 presents the workflow for the three proposed strategies, and all of the three proposed augmentation strategies can be abstracted into at most four components if applicable as follows:
  • Retriever Given input code samples (i.e., vulnerable and clean samples), Retriever is responsible for seeking suitable vulnerable and non-vulnerable pairs from the database and attaching them in the prompt to provide context for vulnerable sample generation if applicable. Injection and Extension strategies employ the Retriever component, while Mutation does not require a Retriever.
  • Formulator For different strategies, we employ their corresponding prompt template. The Formulator instantiates the corresponding prompt template by filling input code samples or the retrieved pairs from Retriever if applicable. For the Mutation prompt template, the input comes directly from the vulnerable samples dataset, while it comes in the form of pairs of clean and vulnerable samples from the Retriever for other templates.
  • Generator In the Generator, we use the instantiated templates to prompt the LLM to generate vulnerable code samples. If the response does not contain a code snippet or we face an API error, a retry mechanism is activated to feed the prompt to the LLM up to 3 times to receive a response that contains a code snippet, otherwise, we move on to the next prompt.
  • Verifier The point of the Verifier is to have control over the quality of the generated code as there’s no guarantee that the LLM produces acceptable code. We use Joern’s (QwietAI, [n. d.]), a fuzzy C language parser, which is based on Antlr’s (Parr and Quong, 1995) C Parser, to filter out generated codes that contain severe syntax errors. The rationale behind using a fuzzy parser instead of a strict parser is that in data augmentation for various machine learning tasks, the generated data does not need to conform to the original data’s strict standards (DeVries and Taylor, 2017; Zhang et al., 2018; Guo et al., 2019; Li et al., 2022; Dong et al., 2023). Note that our goal in this study is to augment existing vulnerable datasets to help models capture the vulnerability patterns and generalize better, rather than generating high-quality vulnerable samples. Completely correct code is not necessary for training models (Wang et al., 2024) and data with subtle noise helps with the generalization of models (Bishop, 1995; Frenay and Verleysen, 2014; Hickey, 1996). Nevertheless, we acknowledge that more sophisticated methods can be used for the verification phase, but starting from a simple module (i.e. a parser) was necessary for evaluating whether using LLMs for vulnerability detection is even feasible. As a results, the fuzzy C Parser filters out 2%-13% of the generated data.
Note that all of our three strategies share the same Generator and Verifier components, while different strategies have slightly different designs for Retriever and Formulator components. We discuss the details of the Retriever and Formulator components for each strategy in Section 3.1.
3.1. Augmentation Strategies
In this section, we break down the details of Mutation, Injection, and Extension strategies. Specifically, we introduce the prompt template and Retriever for each strategy.
3.1.1. Mutation
The majority of previous works on program augmentation (Yu et al., 2022; Orvalho et al., 2022; Mao et al., 2024; Bui et al., 2021; Dong et al., 2023; Chakraborty et al., 2022), rely on code transforms that do not change the flow, semantics, and syntactical correctness of the program using program analysis. Variable name changing, replacing for with while loops and vice versa, and adding dead code are examples of this. However, program analysis is very time-consuming, and selecting the locations for transformation is challenging. Previous studies usually select the types and locations of transformation randomly without any program comprehension (Yu et al., 2022; Orvalho et al., 2022; Mao et al., 2024; Bui et al., 2021; Dong et al., 2023; Chakraborty et al., 2022). In Mutation strategy, we aim to augment vulnerable code samples by utilizing the program comprehension capability of LLM to mutate existing vulnerable code samples and let the LLM choose both the type of transformation and the potential statements to transform. In this way, we get around the program analysis and rely on LLM’s creativity to generate more suitable and diverse items. We design the prompt template as shown in Figure 2.
Prompt Template: Mutation Here’s a code snippet including a function. Except for the important lines mentioned below, mutate the rest of the code snippet so it is different from the original while keeping the original semantics. To do so you can use one or more of the following rules:
Rules: 1- Replace the local variables’ identifiers with new non-repeated identifiers
2- Replace the for statement with a semantic-equivalent while statement, and vice versa
3- {…remaining_rules_are_hidden_to_save_space}
Code Snippet: {input_vulnerable_sample}
The following lines are important and should be included in the final result, but they can still be changed using only the first 5 rules, the rest may be changed using any of the rules or can even be removed if they have no relation to these lines: (Lines are separated via /~/) {input_vulnerable_lines}
Put the generated code inside “‘C “‘.
(Do not include comments)
Figure 2. The mutation prompt template.
We instruct the LLM to use one of the 18 program transformation rules by following a recent study (Yu et al., 2022) and mutating vulnerable code samples while using at least one of the rules. When designing the template, we prioritize the following points. First, we aim to generate diverse code. For this, we mention all of the rules in a single prompt instead of having a prompt per rule to give the LLM the freedom to choose the more suitable rules and apply them to the vulnerable code snippet. Second, we seek to keep the existing vulnerability unchanged in the code. As some of the rules might change the execution trace of the code (e.g. transforming a “switch-case” to an “if-else” statement) and potentially change the status of the vulnerability, we instruct the LLM to use such rules only on the lines that are not important. Important lines are essentially the vulnerable lines of the code snippet. We refer to vulnerable lines as important lines in the prompt to prevent unwanted changes in the resulting code. However, these lines can still be transformed with the rules (i.e., the first five rules in the template) that do not change the execution trace without changing the vulnerable state of the code snippet.
As presented in Figure 1, the overall workflow for Mutation is straightforward. To augment $N$ vulnerable samples, $N$ vulnerable samples are randomly sampled from the input dataset, and then are directly fed into the Formulator to instantiate the template, as a result, $N$ prompts are instantiated by filling the templates. The prompts are then fed to the Generator to generate $N$ vulnerable samples. It’s worth noting that on average 3% of the generated samples will be filtered out in the Verifier component and so if one desires to end up with at least $N$ samples, a higher target should be selected in the generator phase, or else they should redo the generation after the verification to create more items to reach the target.
3.1.2. Injection
Similar to recent works (Nong et al., 2024, 2023, 2022), we also focus on injecting vulnerable code segments into a clean sample as our main strategy, but we aim to cover all types of vulnerabilities and not just single statement vulnerabilities. More specifically, we instruct the LLM to inject the logic of the vulnerable sample into a clean sample by prioritizing the injection of vulnerable segments. Using an LLM for injecting the vulnerable segments into a clean item gives the freedom to LLM to identify the best location. We present our Injection template prompt as shown in Figure 3.
Prompt Template: Injection Here are two code snippets specified below, modify code snippet 2 in a way that it includes the logic of code snippet 1:
Code Snippet 1: {input_vulnerable_sample}
Code Snippet 2: {input_clean_sample}
Note that the following lines from the first code snippet have a high priority to be in the final modified code:
Lines separated by /~/: {input_vulnerable_lines}
Put the generated code inside “‘C “‘.
(Do not include comments)
Figure 3. The injection prompt template.
The templates contain two or three placeholders to fill, namely input_vulnerable_sample, input_clean_sample, and input_vulnerable_lines. input_clean_sample is the input clean samples where we aim to inject vulnerability and is not present in the Mutation strategy. input_vulnerable_sample is a vulnerable code example to be retrieved that is similar to the clean input sample. We retrieve similar vulnerable code because we believe that the logic from a similar vulnerable code could be integrated into the clean code more easily and naturally. input_vulnerable_lines provides meta-information indicating the vulnerable lines in the retrieved vulnerable example. We instruct the LLM to prioritize including these vulnerable lines since our goal is to generate vulnerable code.
To fill the template, we need to construct a dataset of pairs of clean samples and their corresponding retrieved similar vulnerable code samples (i.e., clean-vul pairs). Note that we choose to provide a clean code sample as the input and search for similar vulnerable samples, rather than providing a vulnerable code sample and searching for similar non-vulnerable samples. This strategy is faster and more efficient due to the smaller number of vulnerable samples, which reduces the dataset size for the retrieval process.
Algorithm 1 outlines our method for constructing a dataset of clean-vul pairs through a retrieval process. Given a dataset containing both vulnerable samples ($V$) and clean samples ($C$), our goal is to retrieve $N$ clean-vul pairs. A straightforward approach would be to retrieve the most similar vulnerable samples for each clean sample, sort them in descending order of similarity, and select the top $N$ pairs. However, focusing solely on similarity reduces the diversity of the retrieved samples, which is counterproductive for data augmentation. Previous studies have shown that higher diversity in the training dataset improves the generalizability of deep learning models (Song et al., 2016; Gong et al., 2019). Therefore, to enhance diversity, we incorporate a clustering phase to group the vulnerable samples into $G$ clusters, ensuring that samples from all clusters are considered for selection.
First, we cluster the vulnerable samples into $G$ groups (line 2). To do this, we use CodeBERT (Feng et al., 2020) to embed each code sample into a vector, specifically utilizing the [CLS] token’s embedding for each sample. We then apply KMeans for clustering, using cosine similarity to measure the similarity of each pair of samples.
Following clustering, we create an index on each cluster using Lucene to facilitate efficient search (lines 7-10). For each clean sample, we search for the most similar vulnerable sample within each cluster using the cluster index, and store the resulting clean-vul pairs for further processing (lines 12-17). To find similar vulnerable samples for each clean sample, we use the BM-25 algorithm (Robertson and Zaragoza, 2009), as previous studies have shown that there is no significant difference between sparse methods like BM-25 and dense methods like CodeBERT when leveraging RAG (Chen et al., 2024; Gao et al., 2023; Nashid et al., 2023). We use Lucene’s implementation of BM-25 for this study.
Once the similarity score is calculated for each pair of clean-vul, we sort the results within each cluster based on the similarity score (lines 19-21). Then, we iteratively select the top pair from each cluster, ensuring a diverse yet relevant selection (lines 25-32). We also sort the groups based on size and start from the largest group ensuring a higher coverage for higher values of $G$ (line 23) by following previous study (Ma et al., 2024). We set $G=5$ in our study.
Input: $V$: Vulnerable samples
Input: $C$: Clean samples
Input: $N$: Number of pairs to sample
Input: $G$: Number of groups into which the vulnerable samples will be clustered
Output: $D$: Dataset with $N$ Clean-Vulnerable pairs
1
[…]
6
[…]
8// Index vulnerable samples of each cluster separately
[…]
10
[…]
13 end for
14
15// each clean item is matched with the most similar vulnerable item of each cluster
16 foreach $\textit{clean}\in C$ do
[…]
20       end for
21
22 end foreach
23// Sort each cluster’s pairs based on their similarity score
[…]
25       $sort\_by\_score(\textit{clustered\_pairs}[cluster\_id],"desc")$
26 end for
27
28// sort the arrays based on their size in descending order
[…]
31 while size($D$) <$N$ do
[…]
33      $\textit{index}\leftarrow i//\textit{G}$
[…]
35             $D.\text{append}(\textit{pairs}[\textit{index}])$
36       end if
37      $i\leftarrow i+1$
38 end while
39
return $D$
Algorithm 1 (Clean, Vulnerable) Pair Retrieval
After the clean-vul pairs are retrieved, those pairs will be fed into Formulator to instantiate the Injection template and go through the rest of the components in the workflow to generate $N$ vulnerable new samples similar to Mutation.
3.1.3. Extension
While injecting vulnerable code segments into clean code is the defacto yet effective way of augmenting vulnerabilities, extending an already vulnerable code snippet by adding additional logic can also generate a new vulnerable sample, which enriches the context where vulnerable code happens. Adding certain parts from a clean sample to an already vulnerable sample has a higher chance of keeping the context of the original vulnerable code intact compared to injecting vulnerable segments into a clean item. By extending an already vulnerable sample, only a small section will be irrelevant to the vulnerable part of the code which is the important part for the models, while injecting vulnerable segments into a clean sample results in a code snippet where the vulnerable section has little connection to the original context and gives itself away. Therefore, we aim to explore it as an alternative to vulnerability injection. We refer to this strategy as Extension. We present the prompt template for Extension in Figure 4.
Prompt Template: Extension Here are two code snippets 1 and 2, each including a function. Add some parts of the logic of 1 to 2 in a way that the result would include important lines of 2 with some parts from 1. You can add variables to 2 to produce a more correct code.
Code Snippet 1: {input_clean_sample}
Code Snippet 2: {input_vulnerable_sample}
The following lines from 2 are important and should be included in the final result, the rest may be changed or even removed if they have no relation to these lines: (Lines are separated via /~/) {input_vulnerable_lines}
Put the generated code inside “‘C “‘ and note that the final result should be a function that takes all the input args of 2 and more if required.
(Do not include comments)
Figure 4. The extension prompt template.
Similar to the Injection strategy, we instruct the LLM to add sections of the logic of the clean sample to the vulnerable sample, while enforcing a no-change policy on the vulnerable lines. The only difference is that rather than injecting vulnerable samples into clean samples, the Extension strategy does the reverse. Therefore, we can exactly reuse the workflow of the Injection strategy by modifying the Retriever, where we aim to retrieve similar clean samples for given vulnerable samples and keep all other components the same. Note that we also tried instructing the LLM to extend vulnerable samples freely without providing clean samples, which offers the LLMs more freedom in extending the vulnerable samples. However, it does not work well, and tends to hallucinate in most cases. Specifically, The LLM usually leaves the vulnerable sample unchanged or behaves like a mutator (e.g., mutating variable names or adding dead code). The LLMs perform poorly at generating code when they are not given clear instructions, hence we designed our strategies in a way that we simply ask the LLM to follow clear instructions (i.e. mix two code pieces) without mentioning the word vulnerability and without asking the LLM anything vague or too general. This prevents the LLM from focusing on the vulnerability aspect of the codes and encourages following the given instructions, rather than hallucinating.
4. Experimental Setting
In this section, we present our research questions (RQs), the datasets, DLVD models, LLMs, evaluation metrics, our analysis approach for each of the RQs, and implementation details.
4.1. Research Questions
We evaluate VulScribeR from different aspects to answer the following research questions.
  • RQ1: How effective is VulScribeR compared to SOTA approaches?
  • RQ2: How does RAG contribute to VulScribeR?
  • RQ3: How does the quantity of the generated samples impact the effectiveness of vulnerability detection models?
In RQ1, we evaluate the effectiveness of VulScribeR for improving the performance of DLVD models by augmenting vulnerable data, by comparing it to current SOTA approaches. In RQ2, we aim to investigate the effectiveness of our design for the Retriever component. In RQ3, we explore how the number of generated samples by our approach and SOTA methods affect the performance of DLVD models.
Table 1. The Studied Datasets
Vul. Samples Clean Samples Total Ratio
Devign 10768 12024 22792 1:1.1
Reveal 2240 20494 22734 1:9.1
$\textbf{BigVul}_{\textit{Train}}$ 8783 142125 150908 1:16.2
$\textbf{BigVul}_{\textit{Validation}}$ 1038 17826 18864 1:17.2
$\textbf{BigVul}_{\textit{Test}}$ 1079 17785 18864 1:16.5
4.2. Datasets
In this study, we evaluate our approach on three different widely used vulnerability detection datasets (Nong et al., 2023, 2024; Yang et al., 2023; Fu and Tantithamthavorn, 2022; Zhou et al., 2019; Chakraborty et al., 2020), namely Devign (Zhou et al., 2019), Reveal (Chakraborty et al., 2020) and BigVul (Fan et al., 2020). Table 1 shows the details of these datasets. All of these datasets include C/C++ functions gathered from real-world projects.
4.2.1. Devign
Following previous studies (Nong et al., 2024, 2023) we use Devign as our primary training dataset as it has a better balance between the two categories, namely the vulnerable and clean samples (i.e. non-vulnerable samples) by following a previous study, VulGen (Nong et al., 2023). We also use the clean samples of this dataset as the clean inputs for the Retriever component.
4.2.2. Bigvul
We use all three sets (i.e., training, validation, and testing set) of BigVul (Fan et al., 2020) for different purposes. We use Bigvul’s training set as the source for collecting vulnerable samples to feed it to the Retriever as the vulnerable samples. However, not all of these items can be used as they lack the vulnerable lines metadata that are required for the Injection and Extension prompt templates. Hence, we only use the 6,610 vulnerable samples from BigVuls’s training set that include the metadata of vulnerable lines. We use the testing set of BigVul as one of the testing sets in our study.
4.2.3. Reveal
Similar to VulGen (Nong et al., 2023), we use Reveal as another testing set in our study. Reveal is a slightly larger dataset than BigVul’s testing set and has twice the number of vulnerable samples.
4.3. Deep learning-based vulnerability detection (DLVD) Models
There are two families of DLVD models, token-based and graph-based. To represent both families, we select a SOTA token-based model and two SOTA graph-based models, namely LineVul (Fu and Tantithamthavorn, 2022), Devign (Zhou et al., 2019), and Reveal (Chakraborty et al., 2020) models. For Devign and Reveal models, we use the same setting as VulGen (Nong et al., 2023) and train both of these models 5 times with random seed values, test them and report the results with the highest F1-score achieved. For Linevul, we follow the settings of VGX (Nong et al., 2024), train for 10 epochs, and select the checkpoint that achieved the highest F1 score on Bigvul’s validation set. We specifically train and test the models with different datasets, following the settings of VulGen (Nong et al., 2023) and VGX (Nong et al., 2024), to ensure there is no information leakage and the results can be trusted to apply to real-world unseen data.
4.4. Evaluation metrics
DLVD is inherently a classification task as a category is assigned to the code snippet. Hence, to evaluate the effectiveness of the studied methods, we use the common classification metrics (Powers, 2011), in line with previous related studies (Yang et al., 2023; Powers, 2011; Li et al., 2018; Wang et al., 2023b; Rajbahadur et al., 2022, 2021; Nong et al., 2023, 2024), namely Precision, Recall, and F1-Score.
4.5. Base LLMs
In our study, we use two different LLMs to generate vulnerable code, namely ChatGPT3.5 (OpenAI, 2022), a commercial general-purpose LLM, and CodeQwen1.5 (Team, 2024), an open-source LLM specialized for code tasks.
For ChatGPT, We use the “gpt-3.5-turbo”, variant of OpenAI’s ChatGPT (OpenAI, 2022) which has shown great potential in previous studies (Wang et al., 2023a; Gao et al., 2023; Chen et al., 2024; Sobania et al., 2023). We set the temperature hyper-parameter to 0.5 since we wish to reduce the randomness of the response to a degree that ChatGPT would follow our instructions, but would not limit its freedom of creativity. We left all the other hyper-parameters to their default values. For CodeQwen, we use the 7B variant of CodeQwen1.5 (Team, 2024), “CodeQwen1.5-7B-Chat”, which is a variant of the general purpose Qwen (Bai et al., 2023). CodeQwen1.5, similar to ChatGPT, is a decoder-only transformer-based model with the difference that it was pre-trained on code data, and supports a larger context length. We did not change any of the hyper-parameters for this model and used the defaults, except for the number of new tokens that we set to ChatGPT’s default (i.e. 4096 tokens).
4.6. Baselines
To compare our studies with previous work and demonstrate the effectiveness of VulScribeR we consider the following baselines:
  • NoAug: This is the most basic of our baselines, where we use the original dataset for training without introducing any changes.
  • ROS: As Yang et al. (Yang et al., 2023) show that Random Oversampling (ROS) can have a considerable effect on improving the performance of DLVD, thus we also ROS as another baseline.
  • VulGen Vulgen (Nong et al., 2023) is a SOTA vulnerability generation method that mines single-statement vulnerabilities and uses a transformer model for locating where a vulnerable pattern should be used for injection. We use this baseline as it can generate a significant amount of data that can be used to improve the performance of DLVD models.
  • VGX: VGX (Nong et al., 2024) is the improved version of VulGen that uses a larger dataset for mining single-statement patterns and employs a more sophisticated localization model. Hence, we use VGX as it is potentially a more powerful method and can generate more diverse high-quality data than VulGen.
4.7. Approach for RQs
4.7.1. Approach of RQ1
To demonstrate the effectiveness of VulScribeR and examine which of the proposed strategies is superior to others, we first generate 5K vulnerable samples by employing different approaches (i.e., VulScribeR and baselines). More specifically, to sample 5K items with any of our strategies, we first aim to generate up to 6K samples with the Generator which are then filtered by the Verifier to yield more than 5K generated samples, from which we sample randomly to get 5K items. For VGX and VulGen, we use the generated results in VGX (Nong et al., 2024) and sample 5K from their results randomly. For ROS, we randomly up-sample 5K samples from the vulnerable samples in the Devign dataset.
We then train the three DLVD models using the augmented Devign dataset. To augment the Devign dataset, we add the 5K generated vulnerable samples to it and add the proportional number of clean samples from the vulnerability benchmark proposed in (Lin et al., 2019) to maintain the original ratio of the dataset, by following previous studies (Nong et al., 2023, 2024). We use BigVul’s testing set and Reveal dataset to evaluate the performance of all of these models. We assess the performance of a vulnerable data augmentation approach by assessing the performance of a specific DLVD with the augmented dataset of the specified strategy. Hence, each strategy will be tested in 12 instances (i.e. 3 DLVD Models * 2 Testing Datasets * 2 LLMs).
4.7.2. Approach of RQ2
In this RQ, we investigate the usefulness of our proposed RAG for the Injection and Extension strategies. Specifically, we eliminate the Retriever component, and instead, we match the clean and vulnerable samples randomly to examine the effect of RAG (i.e. Injection w/o Retriever and Extension w/o Retriever).
Moreover, we also conduct an ablation study on the clustering phase of the Retriever component to examine the effect of diversity on the quality of the generated samples. To do so, we remove the clustering phase from the Retriever component and search for similar samples for each clean input item within all the vulnerable items instead of within each cluster. Then we take the top 5 retrieved items to end up with a similar number of $clean-vul$ pairs. In the sampling phase, we sort all the pairs based on their relevance score and start from the top. We denote the Injection strategy without clustering phase as Injection w/o Clustering and the Extension strategy without cluster as Extension w/o Clustering.
Similar to RQ1, we generate 5k vulnerable samples using the studied approaches and compare their effectiveness by assessing the DLVD models that are trained on the augmented datasets.
4.7.3. Approach of RQ3
Lastly, we investigate how the number of augmented samples impacts the performance of DLVD models, and how much performance can be gained by providing more data.
To see the impact and find out whether VulScribeR can be used for large-scale vulnerability augmentation, we use the Injection strategy, our best strategy as depicted in Section 5, and generate up to 16.5K samples so that after the filtering we end up with a set of at least 15K vulnerable samples which is about 40% higher than the number of vulnerable samples in the original dataset. Then we repeat the experiments for the Injection using both LLMs. We sample 10K and 15K vulnerable items while adding the proportional number of clean items to maintain the ratio. We do the same for VGX (Nong et al., 2024), VulGen (Nong et al., 2023), and ROS.
4.8. Implementation details
We use four 24GB Nvidia RTX 3090s for loading and prompting CodeQwen1.5, training, and testing LineVul, Devign, and Reveal models. We conducted a total of 616 experiments for all our RQs which cost more than 1500 GPU hours solely for training and testing DLVD models (excluding the time spent generating the samples).
5. Results
5.1. RQ1: How effective is VulScribeR compared to SOTA approaches?
Table 2. Comparison of VulScribeR’s strategies with baselines using ChatGPT and CodeQwen when augmenting 5K Samples. The cells with larger values (better performance) compared to NoAug are highlighted darker.
ChatGPT 3.5 Turbo CodeQwen1.5-7B-Chat
Strategy Devign Reveal Linevul Devign Reveal Linevul
P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
Reveal NoAug 10.59 47.23 17.30 12.90 65.14 21.54 5.57 16.98 8.39 10.59 47.23 17.30 12.90 65.14 21.54 5.57 16.98 8.39
VulGen 10.45 57.90 17.70 12.46 64.60 20.90 5.31 98.09 10.08 10.45 57.90 17.70 12.46 64.60 20.90 5.31 98.09 10.08
VGX 11.02 59.35 18.58 11.69 65.62 19.84 9.51 87.24 11.47 11.02 59.35 18.58 11.69 65.62 19.84 9.51 87.24 11.47
ROS 13.28 39.93 19.93 14.35 54.52 22.72 9.88 22.71 13.77 13.28 39.93 19.93 14.35 54.52 22.72 9.88 22.71 13.77
Mutation 10.98 53.44 18.22 12.23 68.46 20.75 7.80 22.62 11.60 10.96 61.64 18.60 15.01 57.96 23.85 9.48 30.43 14.45
Injection 13.44 53.26 21.46 15.43 61.46 24.66 11.57 26.43 16.10 13.67 54.28 21.83 14.23 64.48 23.32 9.99 25.79 14.40
Extension 14.97 37.21 21.35 14.76 50.00 22.80 18.01 12.81 14.97 15.20 38.18 21.75 14.82 52.71 23.13 20.65 12.62 15.67
Bigvul NoAug 6.27 43.24 10.95 7.28 77.27 13.31 7.90 29.19 12.43 6.27 43.24 10.95 7.28 77.27 13.31 7.90 29.19 12.43
VulGen 6.49 62.32 11.75 6.36 70.59 11.66 8.51 22.150 12.29 6.49 62.32 11.75 6.36 70.59 11.66 8.51 22.15 12.29
VGX 6.06 59.30 10.99 6.89 73.29 12.60 5.35 98.82 10.15 6.06 59.30 10.99 6.89 73.29 12.60 5.35 98.82 10.15
ROS 7.54 25.60 11.65 7.96 33.39 12.86 10.91 13.16 11.93 7.54 25.60 11.65 7.96 33.39 12.86 10.91 13.16 11.93
Mutation 7.80 63.91 13.90 7.63 67.73 13.72 6.78 74.14 12.42 7.56 57.23 13.36 9.37 53.26 15.94 6.65 78.31 12.25
Injection 8.79 29.41 13.53 11.64 38.00 17.82 11.43 18.81 14.22 10.33 31.80 15.59 9.91 38.16 15.74 7.85 30.49 12.49
Extension 10.60 17.33 13.16 11.13 34.02 16.77 19.03 11.96 14.68 9.89 31.48 15.05 10.22 33.86 15.70 15.74 11.86 13.53
Extension and Injection outperforms Mutation. Injection slightly outperforms Extension. Table 2 presents the results of VulScribeR and baselines across different experimental instances. As observed, Extension and Injection beat Mutation in most (9 out of 12) instances. Extension and Injection perform close to each other, while Injection appears to have the advantage by a tiny margin in most (9 out of 12) settings. In terms of F1-score, Injection achieves an F1-score of 17.60%, outperforming Extension (F1-score 17.38%) and Mutation (F1-score 15.75%) by 0.96% and 12.44% on average, respectively. One possible reason is that Injection and Extension enrich the context where vulnerabilities could occur while Mutation does not alter the semantics of vulnerable code. Mutation does not enrich the context of the semantics of the original vulnerable code.
Figure 5. Entropy of the augmented vulnerable datasets by different approaches. Higher indicates more diversity.
All our proposed LLM-based augmentation strategies outperform the baselines. Typically, Injection and Extension always outperform baselines in all instances. Injection outperforms NoAug, Vulgen, VGX, and ROS by 30.80%, 27.48%, 27.93%, and 15.41% on average F1-score, respectively. As shown in Table 2, Injection and Extension strategies outperform baselines (i.e., NoAug, VulGen, VGX, ROS) in all of the experimental instances in terms of F1-score. More specifically, Extension beats the baselines: NoAug, Vulgen, VGX, and ROS by 29.68%, 26.27%, 26.90%, and 14.35% on average F1-score, while the Injection strategy beats the baselines: NoAug, Vulgen, VGX, and ROS by 30.80%, 27.48%, 27.93%, and 15.41% on average F1-score respectively.
As observed, Mutation beats most baselines by a large margin. However, Mutation does not have a dominating advantage over ROS, and ROS beats Mutation in some (4 out of 12) instances. It is worth noting that ROS beats the baseline, Vulgen, and VGX by 13.96%, 10.72%, and 11.21% in terms of F1-score on average.
As discussed in Section 2, SOTA approaches like VGX and VulGen only focus on single-statement that limits the diversity of generated vulnerabilities. However, our approaches do not have this limitation and are able to generate more diverse vulnerable samples. To examine this, we measure the diversity of the augmented vulnerable samples by our approaches and VGX and VulGen. By following previous studies, we first apply principle component analysis (PCA) on CodeBERT’s embeddings of the generated vulnerable samples and reduce the dimension to three. Next, we calculate the histogram using 10 bins, from which we calculate the entropy of the vectors. Higher entropy values indicate greater diversity among the vulnerable samples, while lower values suggest a more concentrated distribution with similar samples. As depicted in Figure 5 the entropy for Mutation, Injection, and Extension are 4.79, 4.5, and 4.62 for ChatGPT, and 4.66, 4.45, 4.42 for CodeQwen, while for Vulgen and VGX, it is 4.42 and 4.31 respectively. The results demonstrate that our approaches have a better potential for generating more diverse samples as we beat them while using only one of the datasets they used for vulnerability mining. VulScribeR produces more diverse vulnerable samples compared to SOTA approaches VXG and VulGen.
Lastly, it is worth noting that the results on both of the LLMs perform very close to each other. CodeQwen1.5-7B-Chat slightly outperforms ChatGPT3.5 Turbo by a tiny margin (i.e. 1.49%) averaged across all three strategies, yet one cannot conclude that CodeQwen is a better LLM for vulnerability generation as we did not explore the hyper-parameters for the optimum setting for each of the LLMs. However, we can conclude that our strategies work with similar LLM to ChatGPT and CodeQwen (e.g. GPT4 (Achiam et al., 2023) and DeepSeek-Coder (Guo et al., 2024)) and do not require an LLM that is trained specifically on code.
Both Injection and Extension outperforms the baselines and the Mutation by a large margin, while Injection provides a slightly higher performance compared to the Extension. For instance, Injection outperforms NoAug, Vulgen, VGX, and ROS by 30.80%, 27.48%, 27.93%, and 15.41% on average F1-score.
5.2. RQ2: How does RAG contribute to VulScribeR?
Table 3. Results of Ablation Studies on the Retriever and Clustering for Injection.
ChatGPT 3.5 Turbo CodeQwen1.5-7B-Chat
Strategy Devign Reveal Linevul Devign Reveal Linevul
P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
Reveal w/o Retriever 13.91 44.33 21.18 14.17 64.29 23.22 8.11 32.15 12.95 11.94 57.12 19.75 12.52 72.74 21.37 13.00 32.06 18.50
w/o Clustering 14.75 36.37 20.99 15.59 59.89 24.74 9.92 24.61 14.14 13.86 43.37 21.01 14.92 52.17 23.20 12.23 24.61 16.34
Injection 13.44 53.26 21.46 15.43 61.46 24.66 11.57 26.43 16.10 13.67 54.28 21.83 14.23 64.48 23.32 9.99 25.79 14.40
Bigvul w/o Retriever 8.47 36.09 13.72 8.96 65.50 15.76 6.67 66.54 12.12 8.87 53.74 15.23 8.88 64.39 15.60 7.09 61.72 12.72
w/o Clustering 8.53 33.23 13.57 11.19 32.75 16.68 9.56 16.68 12.16 9.77 32.75 15.05 9.49 48.65 15.88 14.61 21.78 17.49
Injection 8.79 29.41 13.53 11.64 38.00 17.82 11.43 18.81 14.22 10.33 31.80 15.59 9.91 38.16 15.74 7.85 30.49 12.49
Table 4. Results of Ablation Studies on the Retriever and Clustering for Extension.
ChatGPT 3.5 Turbo CodeQwen1.5-7B-Chat
Strategy Devign Reveal Linevul Devign Reveal Linevul
P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
Reveal w/o Retriever 13.96 35.59 20.05 14.00 49.94 21.87 8.63 12.20 12.20 11.55 67.73 19.74 13.21 65.26 21.98 11.52 32.15 16.96
w/o Clustering 15.14 36.01 21.32 14.34 63.63 23.41 7.82 48.38 13.46 15.68 36.85 22.00 14.93 50.84 23.09 9.77 21.34 13.40
Extention 14.97 37.21 21.35 14.76 50.00 22.80 18.01 12.81 14.97 15.20 38.18 21.75 14.82 52.71 23.13 20.65 12.62 15.67
Bigvul w/o Retriever 7.88 15.42 10.43 8.62 25.12 12.83 7.13 72.75 12.99 7.66 67.57 13.77 8.91 62.48 15.60 6.89 67.66 12.50
w/o Clustering 10.78 33.70 16.33 11.13 22.58 14.91 8.20 17.98 11.26 12.52 30.21 17.70 9.83 39.75 15.76 9.38 26.41 13.85
Extention 10.60 17.33 13.16 11.13 34.02 16.77 19.03 11.96 14.68 9.89 31.48 15.05 10.22 33.86 15.70 15.74 11.86 13.53
For Injection, the clustering phase of the Retriever adds 0.49% to the F1-score averaged across all settings compared to w/o clustering, and w/o clustering improves the average f1-score by 6.34% compared to w/o Retriever. As shown in Table 3, both similarity and diversity affect the effectiveness of the models positively with similarity having a more significant impact. When comparing Injection with Injection w/o Retriever, we observe that the complete Retriever component improves the effectiveness of the Injection by 4.99% on the average F1-score and in 8 out of 12 instances. Specifically, when we compare Injection with w/o the clustering phase, we see that the clustering phase enhances the effectiveness of RAG by 0.49% on the average F1-score and in 8 out of 12 instances, and by comparing w/o clustering with w/o Retriever, we see that RAG enhances the effectiveness of RAG by 6.34% on the average F1-score and in 8 out of 12 instances.
For Extension strategy, the clustering phase of The Retriever provides 2.54% of improvement to the F1-score averaged across all settings compared to w/o clustering, and w/o clustering improves the average f1-score by 5.61% compared to w/o Retriever. As shown in Table 4, both similarity and diversity affect the effectiveness of the models positively with similarity having a more significant impact. Overall, our complete Retriever module improves the effectiveness of the Extension strategy by 10.77% on the average F1-score and beats the other settings in 6 out of 12 settings. Specifically, by comparing Extension with w/o clustering we see that the clustering phase enhances the effectiveness of RAG by by 2.54% on the average F1-score and in 7 out of 12 instances, and by comparing w/o clustering with w/o Retriever, we see that RAG enhances the effectiveness of RAG by by 5.61% on the average F1-score and in 10 out of 12 instances.
Retriever component makes significant contribution for Injection and Extension. Extension gains more than twice the increase from the Retriever component, implying that it is more sensitive to both the similarity and diversity of the retrieved pairs.
5.3. RQ3: How does the quantity of the generated samples impact the effectiveness of vulnerability detection models?
Table 5. Impact of Generated Samples on Improving DVLD models’ Performance at 5K, 10K, and 15K
ChatGPT 3.5 Turbo CodeQwen1.5-7B-Chat
Strategy Devign Reveal Linevul Devign Reveal Linevul
P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
Reveal NoAug 10.59 47.23 17.30 12.90 65.14 21.54 5.57 16.98 8.39 10.59 47.23 17.30 12.90 65.14 21.54 5.57 16.98 8.39
VulGen 5K 10.45 57.90 17.70 12.46 64.60 20.90 5.31 98.09 10.08 10.45 57.90 17.70 12.46 64.60 20.90 5.31 98.09 10.08
VulGen 10K 10.21 46.20 16.72 11.31 66.77 19.34 5.37 100.00 10.20 10.21 46.20 16.72 11.31 66.77 19.34 5.37 100.00 10.20
Vulgen 15K 10.30 55.25 17.37 11.23 75.51 19.55 5.35 99.09 10.15 10.30 55.25 17.37 11.23 75.51 19.55 5.35 99.09 10.15
VGX 5K 11.02 59.35 18.58 11.69 65.62 19.84 9.51 87.24 11.47 11.02 59.35 18.58 11.69 65.62 19.84 9.51 87.24 11.47
VGX 10K 10.31 57.90 17.50 11.53 56.45 19.15 5.24 95.91 9.93 10.31 57.90 17.50 11.53 56.45 19.15 5.24 95.91 9.93
VGX 15K 9.01 34.68 14.30 11.95 63.75 20.13 5.38 99.91 10.21 9.01 34.68 14.30 11.95 63.75 20.13 5.38 99.91 10.21
ROS 5K 13.28 39.93 19.93 14.35 54.52 22.72 9.88 22.71 13.77 13.28 39.93 19.93 14.35 54.52 22.72 9.88 22.71 13.77
ROS 10K 12.92 51.81 20.68 14.00 46.20 21.49 9.39 30.06 14.31 12.92 51.81 20.68 14.00 46.20 21.49 9.39 30.06 14.31
ROS 15K 13.25 47.29 20.71 13.77 43.85 20.96 8.18 19.71 11.56 13.25 47.29 20.71 13.77 43.85 20.96 8.18 19.71 11.56
Injection 5K 13.44 53.26 21.46 15.43 61.46 24.66 11.57 26.43 16.10 13.67 54.28 21.83 14.23 64.48 23.32 9.99 25.79 14.40
Injection 10K 17.39 42.70 24.71 16.37 63.21 26.00 11.19 31.97 16.57 15.87 48.31 23.90 14.96 63.03 24.18 15.05 19.53 17.00
Injection 15K 16.94 52.65 25.63 17.00 57.36 26.22 13.02 36.61 19.21 15.60 52.53 24.06 16.27 57.18 25.33 16.01 23.89 19.17
Bigvul NoAug 6.27 43.24 10.95 7.28 77.27 13.31 7.90 29.19 12.43 6.27 43.24 10.95 7.28 77.27 13.31 7.90 29.19 12.43
VulGen 5K 6.49 62.32 11.75 6.36 70.59 11.66 8.51 22.150 12.29 6.49 62.32 11.75 6.36 70.59 11.66 8.51 22.150 12.29
VulGen 10K 6.59 52.62 11.72 6.21 76.79 11.50 12.04 12.70 12.36 6.59 52.62 11.72 6.21 76.79 11.50 12.04 12.70 12.36
Vulgen 15K 6.35 62.96 11.54 6.48 69.95 11.86 8.32 14.74 10.64 6.35 62.96 11.54 6.48 69.95 11.86 8.32 14.74 10.64
VGX 5K 6.06 59.30 10.99 6.89 73.29 12.60 5.35 98.82 10.15 6.06 59.30 10.99 6.89 73.29 12.60 5.35 98.82 10.15
VGX 10K 5.84 61.21 10.65 6.36 63.28 11.56 7.14 10.10 8.37 5.84 61.21 10.65 6.36 63.28 11.56 7.14 10.10 8.37
VGX 15K 6.11 50.40 10.90 6.42 66.14 11.70 9.36 6.49 7.66 6.11 50.40 10.90 6.42 66.14 11.70 9.36 6.49 7.66
ROS 5K 7.54 25.60 11.65 7.96 33.39 12.86 10.91 13.16 11.93 7.54 25.60 11.65 7.96 33.39 12.86 10.91 13.16 11.93
ROS 10K 7.64 32.43 12.37 9.32 33.55 14.59 8.13 18.44 11.28 7.64 32.43 12.37 9.32 33.55 14.59 8.13 18.44 11.28
ROS 15K 9.31 21.14 12.93 8.22 26.87 12.59 8.77 13.07 10.49 9.31 21.14 12.93 8.22 26.87 12.59 8.77 13.07 10.49
Injection 5K 8.79 29.41 13.53 11.64 38.00 17.82 11.43 18.81 14.22 10.33 31.80 15.59 9.91 38.16 15.74 7.85 30.49 12.49
Injection 10K 13.94 17.60 17.60 12.09 43.56 18.92 8.84 33.09 13.96 11.49 30.37 16.67 11.54 41.81 18.08 9.86 19.83 13.17
Injection 15K 12.70 34.66 18.59 11.92 41.81 18.55 11.78 30.15 16.94 12.12 33.86 17.85 11.54 41.81 18.09 9.74 29.19 14.60
Figure 6. The impact of the number of augmented samples on the effectiveness of DLVD models across the studied datasets and LLMs.
Augmenting more vulnerable data by using Injection helps improve the effectiveness of DLVD models. Table 5 and Figure 6 present the performance (in terms of F-1 score) of DLVD models when being trained on a dataset that is augmented with different vulnerable samples generated by Injection and other baselines (e.g., VulGen, VGX, ROS, and NoAug). As we can see, The performance of DLVD models increases as more augmented vulnerable samples are added to the training data in almost all experimental instances (11 out of 12), which indicates that adding more data augmented by Injection provides more useful information for the model to capture vulnerabilities.
In general, VulGen, VGX, and ROS fail to improve the performance of DLVD models by augmenting more vulnerable samples. When looking at the baselines in Figures 6, we notice that Vulgen, VGX, and ROS struggle with improving the performance of DLVD models by augmenting more data. For instance, if we compare the performance of DLVD models when augmenting 5K and 15K vulnerable samples, VGX, VulGen, and ROS decrease the performance in 5 out of 6, 4 out of 6, and 4 out of 6 instances, respectively (Note that the baselines are not dependent on the LLMs, but they are shown for easier comparison). This finding aligns with the results reported in Yang et al. (Yang et al., 2023), indicating that excessive random over-sampling does not improve the effectiveness of DLVD models and can diminish their performance. One possible explanation is that since random over-sampling does not introduce any new information, and excessive over-sampling probably leads to overfitting of the oversampled data in the models. For VGX and VulGen, this is probably due to the nature of their method, which solely focuses on single-statement vulnerabilities. This heavily limits the usefulness of such methods and makes them unfeasible to be used for large-scale data augmentation.
In summary, Injection outperforms all baselines (i.e., VulGen, VGX, and ROS) in all scales, typically when used for large-scale data augmentation. The studied baselines fail to improve the effectiveness of DLVD models by augmenting more than 5K samples. Our LLM-based approach is more feasible for large-scale vulnerable data augmentation.
Augmenting more vulnerable data by using Injection helps improve the effectiveness of DLVD models, while VulGen, VGX, and ROS fail to improve the performance of DLVD models by augmenting more than 5K vulnerable samples. In contrast, our LLM-based approach is more feasible for large-scall vulnerable data augmentation. Explicitly, Injection at 15K beats NoAug, Vulgen at 15K, VGX at 15K, and ROS at 15K by 53.84%, 54.10%, 69.90%, and 40.93%.
6. Discussion
6.1. Cost analysis
Based on our experiments, generating every 1,000 vulnerable samples with one prompt at a time and no concurrency, takes 3.4 hours and costs about \$1.88 with GPT3.5-Turbo and about 9 GPU hours with CodeQwen1.5-7B-Chat on two RTX3090s. This shows the feasibility of VulScribeR as it takes less than \$19 with ChatGPT to augment a dataset like Devign to twice its size and improves the performance of a DLVD model up to 128.95%.
6.2. Threats to Validity
Internal Validity A common threat to the validity of our study is our hyper-parameter settings for DLVD models and LLMs. For DLVD, hyper-parameter tuning is extremely expensive given the scale of our study, hence we followed the setting of previous studies (Nong et al., 2024, 2023). In addition, when comparing different data augmentation approaches we used the same settings for DLVD models, which ensures our comparisons are fair. For LLMs, we mostly used the defaults, and only set the number of new tokens to ChatGPT’s maximum, and lowered ChatGPT’s temperature to 0.5. Better results might be achievable by fine-tuning these hyper-parameters. Another threat is that LLMs might not always follow the prompts as instructed and could produce hallucinations. To alleviate this threat, we add a verifier component to filter out low-quality code, and we use a retry mechanism that prompts the LLM up to three times in case of errors (including an empty code response). More importantly, our goal in this study is to augment existing vulnerable datasets to help models capture the vulnerability patterns and generalize better, rather than generating high-quality vulnerable samples. Completely correct code is not necessary for training models (Wang et al., 2024) and data with subtle noise helps with the generalization of models (Bishop, 1995; Frenay and Verleysen, 2014; Hickey, 1996). We acknowledge that more sophisticated methods can be used for the verification phase, but starting from a simple module (i.e. a parser) was necessary for evaluating whether using LLMs for vulnerability detection is even feasible.
External Validity relates to the generalizability of our findings. Even though we used two different LLMs (ChatGPT 3.5 Turbo and CodeQwen1.5-7B-Chat), evaluated on three commonly used datasets, and covered three SOTA DLVD models from two different families, our findings may not generalize well to real-world scenarios where different models, LLMs, and datasets are used. To encourage future research to investigate more LLMs (including encoder-decoder-based models), datasets, and DLVD models.
7. Conclusion
In this paper, we propose VulScribeR, a novel yet effective framework for augmenting vulnerable samples that uses a customized RAG mechanism to formulate carefully designed prompt templates which are then used for prompting the LLMs. We also use a fuzzy parser as a verification method, to make sure that the generated code doesn’t have severe syntax problems. Our evaluation results show that our method significantly outperforms SOTA methods, and is suitable for large-scale vulnerability augmentation with low cost.
8. Data Availability
We made our replication package publicly available (dat, 2024) to encourage future research on enhancing the performance of LLM-generated vulnerable code and more advanced filtering mechanisms.